# Expertized Caption Auto-Enhancement for Video-Text Retrieval

### v1

@@ Abstract
The burgeoning field of video-text retrieval has witnessed significant advancements with the advent of deep learning. However, the challenge of matching text and video persists due to inadequate textual descriptions of videos. The substantial information gap between the two modalities hinders a comprehensive understanding of videos, resulting in ambiguous retrieval results. While rewriting methods based on large language models have been proposed to broaden text expressions, carefully crafted prompts are essential to ensure the reasonableness and completeness of the rewritten texts. This paper proposes an automatic caption enhancement method that enhances expression quality and mitigates empiricism in augmented captions through self-learning. Additionally, an expertized caption selection mechanism is designed and introduced to customize augmented captions for each video, facilitating video-text matching. Our method is entirely data-driven, which not only dispenses with heavy data collection and computation workload but also improves self-adaptability by circumventing lexicon dependence and introducing personalized matching. The superiority of our method is validated by state-of-the-art results on various benchmarks, specifically achieving Top-1 recall accuracy of 68.5% on MSR-VTT, 68.1% on MSVD, and 62.0% on DiDeMo.
1 Introduction
Video-text understanding has become a new mainstream research with the population of short videos on social networks. In virtue of the learning ability of large models, multi-modal pre-training models [11, 6] are developed and achieve initial success in video-text retrieval tasks. However, a stunning amount of data and computing resources are the essential basis to support the effectiveness of these pre-training models. Thus, it is argued that adapting the image-text pretrained model to video-text tasks is a lighter and more feasible approach [15, 31]. They adapt the visual embeddings to the video field by incorporating temporal information [21, 4] or refining the visual-text matching strategies [8, 31]. But unlike image-text pairs, video captions tend to be overly simplistic and one-sided, describing partial (even a signal) frames of the videos. These one-sided descriptions would be improperly matched to multifaceted video embeddings (Figure 1 (a)), resulting in biased cross-modal retrieval, and a bottleneck in performance improvement arises.
@@ 1 Introduction
Frame sampling is one proper solution to handle the information imbalance issue. For instance, ClipBERT [10] sparsely samples video frames to meet the text descriptions. X-CLIP [16] further cuts text descriptions into words, employing cross-modal matching at both fine-grained (frame and word) and coarse-grained (video and sentence) levels. As video information is sparsely divided and broken, these methods are limited in comprehensively utilizing and exploring deep relations across videos and texts. Another branch to reduce the information gap is data agumentation. Extra data, such as HD-VILA-100M [30] and LF-VILA [21], are collected to expand the receptive field for model adaptation, thereby enriching matching knowledge of the two modalities. To eliminate the heavy data collection workload, later research [27] rewrites captions by lexical analysis, in which attributes are extracted and emphasized to enhance expressive ability in text modality. Large language models (LLMs), such as Chat-GPT and LLaMA [22], are additionally introduced for the rewriting tasks [26, 7]. Although up to 3% retrieval precision increase is achieved with the assistance of supplementary texts, the rewriting quality relies heavily on lexicon engineering [18] and hand-crafted prompts [17] (Figure 2 (a)). Recent works [2, 25] have been mired in the meticulous work of designing better prompts for caption augmentation, limited in human empiricism.
@@
This paper proposes an Expertized Caption auto-Enhancement (ExCae) method, which adaptively generates and personalized select captions for video-text retrieval, releasing extra work of data acquisition and avoiding the trouble of manual prompt design. Specifically, a Caption Self-improvement (CSI) module is designed to derive captions with rich content and diverse perspectives from videos. A Prompt engineer is introduced in the CSI module, which gradually adjusts and improves the query prompts via self-learning. With the enriched expressions on the video-side (video+video-derived captions), cross-modal matching becomes a more challenging task. Thus, we further develop an Expertized Caption Selection (ECS) module consisting of multiple learnable experts to automatically specify appropriate expressions for cross-modal matching. Seeing Figure 1 (b), video representation space dilates with caption augmentation and the cross-modal information gap is reduced after expertized selection. This indicates that the proposed method improves the generalization ability through caption augmentation while enhancing the cross-modal retrieval ability with adaptive screening. As a fully data-driven process, sampling bias and empiricism that may confuse the cross-modal matching are avoided. We further encapsulate ExCae as a plug-in unit, which successfully improves the retrieval performance of current methods, validating ExCae’s flexibility and adaptability.
@@
  • We propose eliminating the grunt workload of data acquisition and prompt design in caption augmentation through self-improvement and expertized selection, thereby mitigating empirical bias in supplemental texts while facilitating cross-modal matching without extra data.
  • We design an automatic cross-modal learning framework for video-text retrieval, in which video expressions are expanded using a Prompt engineer and are matched to texts under the guidance of learnable experts. The Prompt engineer facilitates generating effective and multi-angled captions while the learnable experts personalize the appropriate expression angles for cross-modal matching.
@@ 2 Related Work
Developing and pre-training video foundation models (VFMs) [11, 24] enhances the comprehension of video data, demonstrating efficacy in downstream tasks such as video-text retrieval. However, the effectiveness of VFMs heavily relies on huge amounts of data and computing resources. Much of the current work in video-text retrieval turns to a lighter modeling approach, adapting image-text retrieval models, such as CLIP [20], to video-text tasks. For instance, CLIP4Clip [15] and CLIP2TV [8] utilize the image knowledge encapsulated in CLIP, transforming it into effective video representations through different image aggregation networks. Besides, TS2-Net [14] dynamically adjusts the video token sequence, selecting tokens with richer information to enhance video representation. X-CLIP [16] employs an independent encoder for each frame, using a network that aggregates frames multiple times to generate the final video representation. DRL [23] introduces an additional loss to address the feature redundant issue. However, due to the information imbalance between videos and texts, these methods are stuck in insufficient cross-modal matching.
@@
To reduce the cross-modal information gap, existing works, such as Cap4Video [26] and BIKE [27], enhance video-text matching via caption rewrite. They carefully design prompts, asking LLMs to rephrase the texts with empirically selective concepts, such as special objectives [18] or actions [17]. However, visual details remain deficient in the rewritten captions through these text-to-text literal restatements. More importantly, the hand-drafted lexicon and prompts severely limit the rewrite quality, motivating us to explore a self-learning caption enhancement method. To the best of our knowledge, this paper is a pioneer in investigating automatic caption enhancement to facilitate the comprehension of videos and texts.
3 Method
@@
Unlike traditional text augmentation that rewrites captions on the text data side, our method (ExCae), automatically derives captions from videos by the CSI module and adaptively selects appropriate expressions (ECS module) to encode video for video-text matching (see Figure 2). Video and text embeddings would be aligned with standard cross-modal contrastive loss, as most existing SOTA methods [14, 15]. Both CSI and ECS are self-learning modules, whose details are introduced as follows.
3.1 Caption Self-improvement (CSI)
The objective of the CSI module is to learn the best prompt for caption generation. The main flow is illustrated in Figure 3, which is an iteration process gradually refining the query prompt. Specifically, a Captioner is asked to generate multiple captions to describe the video from different perspectives. We then evaluate the semantics of the generated captions, scoring them to update the prompt candidates’ pool. A Prompt engineer is introduced to rewrite the current best prompt, whose output is used as the latest query prompt of the Captioner. The prompt update iteration would be ended when no new candidates are added to the prompt candidates’ pool. (Pseudo code is provided in the Supplemental Materials.)
@@ 3.1 Caption Self-improvement (CSI)
Denote the query prompt of the Captioner as $p^{t}$, where $t$ records the iteration index. Correspondingly, the generated caption list for a video $v_{i}$ based on $p^{t}$ is written as $c^{t}_{i}$. As literal expressions, $c^{t}_{i}$ is evaluated from a semantic perspective. We advocate for the generation of captions that are semantically akin to the original texts, in order to mitigate the presence of cluttered and vacuous descriptions. However, merely preserving the semantic consistency may lead to generating a set of captions similar to the original text. This is opposite to the original intention of expression enhancement. To promote multi-angle descriptions, a diversity constraint that encourages variety within the caption list is supplemented. Suppose $K$ captions are contained in $c^{t}_{i}$, the caption generation score in the $t$ iteration round can be expressed as
$Score^{t}=\mathbb{E}_{i,k}[sim(\psi(c^{t}_{i,k}),\psi(t_{i}))]+div(\psi(c^{t}_{i}))$ (1)
where $t_{i}$ denotes the original text corresponding to $v_{i}$, $k\in[1,K]$ is the caption index in $c^{t}_{i}$. $\psi(\cdot)$ represents a semantic extractor, implemented by M3-embedding [5]. $sim(\cdot)$ and $div(\cdot)$ are similarity and diversity calculations (See Supplemental Materials for details). If $Score^{t}$ surpasses the historical records, $p^{t}$ would be added to the prompt candidates’ pool, and the best prompt is updated accordingly.
@@
The current prompt engineering normally encodes the query prompts and adjusts the prompt tokens/embeddings through gradient descent [19, 32]. Although this parametric fine-tuning strategy is convenient for optimization, the adjusted embeddings could hardly be precisely re-projected to the text space, losing linguistic meaning and interpretability. Consequently, their generalization in video-text-related applications is limited. Inspired by language rewrites [7], this paper introduces a Prompt engineer to rewrite the query prompt. Take the current best prompt $p^{t}$ as input; the Prompt engineer is asked to rewrite $p^{t}$ with reference to the corresponding video-text pairs $(x,t)$ involved in this iteration. The rewritten prompt is then used as the query prompt for the Captioner in the next iteration, as shown in Figure 3.
Figure 3: Illustration of the CSI moudle
3.2 Expertized Caption Selection (ECS)
With the optimized CSI module, video-text pairs are updated as $((v,\tilde{c}),t)$, where $(v,\tilde{c})$ represents the video data $v$ and its generated captions $\tilde{c}$ using the optimal prompt, and $t$ denotes the original text corresponding to $v$. As displayed in Figure 2 (b), $(v,\tilde{c})$ and $t$ are the video-side and text-side input data, respectively. A straightforward approach for cross-modal representation learning is directly matching the encoded representations of $(v,\tilde{c})$ and $t$. However, the information dilation of video modality brings confounded knowledge while improving generalization, increasing the difficulty of cross-modal matching (see Figure 1).
@@ 3.2 Expertized Caption Selection (ECS)
Table 1: Comparison with SOTA methods on MSR-VTT dataset.
4 Experiment
4.1 Datasets
MSR-VTT [29] is a large-scale dataset containing 10,000 video clips, each of which is annotated with 20 English sentences. Around 29K unique words are contained in all sentences. Following previous work [13], we use 9,000 clips for training and 1,000 clips for testing.
MSVD [3] consists of 1,970 videos with 80K captions. Each video contains around 40 English captions on average. The videos are split by 1,200/100/670 for training, validation and testing, respectively.
DiDeMo [1] contains over 10,000 videos collected from Flickr. Each video is described by natural language, with detailed information on camera movement, temporal transition indicators, and activities. The dataset is split into training, validation and test sets, each containing 8,395, 1,065 and 1,004 videos, respectively.

### v3

@@ Abstract
Video-text retrieval has been stuck in the information mismatch caused by personalized and inadequate textual descriptions of videos. The substantial information gap between the two modalities hinders an effective cross-modal representation alignment, resulting in ambiguous retrieval results. Although text rewriting methods have been proposed to broaden text expressions, the modality gap remains significant, as the text representation space is hardly expanded with insufficient semantic enrichment. Instead, this paper turns to enhancing visual presentation, bridging video expression closer to textual representation via caption generation and thereby facilitating video-text matching. While multimodal large language models (mLLM) have shown a powerful capability to convert video content into text, carefully crafted prompts are essential to ensure the reasonableness and completeness of the generated captions. Therefore, this paper proposes an automatic caption enhancement method that improves expression quality and mitigates empiricism in augmented captions through self-learning. Additionally, an expertized caption selection mechanism is designed and introduced to customize augmented captions for each video, further exploring the utilization potential of caption augmentation. Our method is entirely data-driven, which not only dispenses with heavy data collection and computation workload but also improves self-adaptability by circumventing lexicon dependence and introducing personalized matching. The superiority of our method is validated by state-of-the-art results on various benchmarks, specifically achieving Top-1 recall accuracy of 68.5% on MSR-VTT, 68.1% on MSVD, and 62.0% on DiDeMo. Our code is publicly available at https://github.com/CaryXiang/ECA4VTR.
Index Terms:
Video-Text Retrieval, Multimodal Representation Learning, Prompt Optimization.
I Introduction
Video-text understanding has become a new mainstream research with the population of short videos on social networks. In virtue of the learning ability of large models, multi-modal pre-training models [1, 2] are developed and achieve initial success in video-text retrieval (VTR) tasks. However, a stunning amount of data and computing resources are the essential basis to support the effectiveness of these pre-training models. A lighter and more feasible approach to achieve VTR is adapting the image-text pretrained model to video-text scenarios [3, 4]. In this approach, the visual embeddings are adapted to the video field by introducing temporal information [5, 6] or refining the visual-text matching strategies [7, 4]. But unlike image-text pairs, the textual descriptions of videos tend to be overly simplistic and one-sided, describing partial (even a single) frames of the videos. The improper alignment between one-sided descriptions and multifaceted video embeddings would confuse text understanding (Fig. 1 (a)), resulting in biased cross-modal retrieval, and a bottleneck in performance improvement arises.
@@ I Introduction
Frame sampling is one proper solution to handle the improper alignment issue. For instance, ClipBERT [8] sparsely samples video frames to meet the text descriptions. X-CLIP [9] further cuts text descriptions into words, searching for video frames for fine-grained cross-modal matching. As video information is sparsely divided and broken, these methods are limited in comprehensively utilizing and exploring deep relations across videos and texts. Another branch to promote video-text understanding is data augmentation. Extra data, such as HD-VILA-100M [10] and LF-VILA [5], are collected to expand the receptive field for model adaptation, thereby enriching matching knowledge of the two modalities. To address the substantial data collection burden, subsequent research has explored text rewriting strategies to enhance expressions on the text side. For example, Wu et al. [11] rewrite texts using lexical analysis, in which attributes are extracted to enhance expressive ability in the text modality. She et al. [12] further ameliorate text presentation using syntax-tree, emphasizing key words to assist better VTR. Besides generating textual descriptions, text enhancement is executed in the feature space [13, 14], providing embedding variants for cross-modal alignment. Despite achieving a retrieval precision increase of up to 3% with the aid of supplementary textual expressions, minimal additional semantics are introduced during the text-to-text rephrasing processes, which hinders further reduction of the modality gap. This restricts significant breakthroughs in cross-modal retrieval capability.
Instead, this paper suggests enhancing the representations on the video side, deriving captions from videos to meet the expression modes and semantic information of texts. The rich semantics of videos are preserved, and video-text matching would be simplified due to the reduced modality gap achieved by this directional data augmentation. Although multimodal Large Language Models (mLLMs) have shown a decent ability in video understanding [15], generating appropriate video captions for VTR remains challenging, as it is revealed that the performance of mLLMs is query-sensitive. The writing quality of mLLMs relies heavily on lexicon engineering [16] and hand-crafted prompts [17]. Recent studies [18, 19] have been mired in the meticulous work of designing better prompts for caption augmentation, limited in human empiricism.
@@
Inspired by automatic prompt optimization (APO) [20, 21], this paper proposes an Expertized Caption Auto-Enhancement (ExCae) method, as shown in Fig. 2 (b). ExCae incorporates two main components: Caption Self-Improvement (CSI) and Expertized Caption Selection (ECS) modules, which adaptively generates and selectively personalizes captions for VTR. The two modules provide ExCae with the advantage of alleviating the additional burden of data acquisition and eliminating the complexities associated with manual prompt design. Specifically, in the CSI module, a Captioner implemented by a multimodal LLM (mLLM) is asked to derive captions from videos. We optimize the query prompt of the Captioner to ensure the generated captions contain rich content and diverse perspectives. To overcome the interpretability loss of prompt in existing gradient-based APO, this paper introduces a Prompt Engineer, which gradually improves the query prompts of the Captioner via repetitively inquest and output adjustment. With the enriched expressions on the video side (video+video-derived captions), the ECS module consisting of multiple learnable experts is designed to automatically specify appropriate expressions for cross-modal matching. Seeing Fig. 1 (b), video representation space dilates with caption augmentation, and the cross-modal information gap is reduced after expertized selection. This indicates that the proposed method improves the generalization ability through caption augmentation while enhancing the cross-modal retrieval ability with adaptive screening. As a fully data-driven process, sampling bias and empiricism that may confuse the cross-modal matching are avoided. We further encapsulate ExCae as a plug-in unit, which successfully improves the retrieval performance of current methods, validating ExCae’s flexibility and adaptability.
@@
  • We propose reducing the modality comprehension gap for VTR by enhancing visual presentations via self-improved caption derivation from videos. Video captions are progressively refined, bridging video expression closer to text space while eliminating the grunt workload of data acquisition.
  • We design an automatic video-text learning framework for VTR, in which video expressions are expanded using a Prompt Engineer and are matched with texts under the guidance of learnable experts. The Prompt Engineer facilitates effective and multi-angle caption generation through a self-improvement strategy that mitigates empirical bias in prompt design, while the learnable experts personalize the appropriate expression angles for cross-modal matching.
@@ II Related Work
Developing and pre-training video foundation models (VFMs) [1, 22] enhances the comprehension of video data, demonstrating efficacy in downstream tasks such as video-text retrieval. However, the effectiveness of VFMs heavily relies on huge amounts of data and computing resources. Much of the current work in video-text retrieval turns to a lighter modeling approach, adapting image-text retrieval models, such as CLIP [23], to video-text tasks. For example, CLIP4Clip [3] and CLIP2TV [7] transform CLIP’s embeddings into effective video representations through various frame aggregation strategies. In addition, X-CLIP [9] proposes using an independent encoder for each frame to derive video representations. Video embeddings are further enhanced in TS2-Net [24], which dynamically selects video tokens with richer information. Transfer learning is another appropriate solution to adapt knowledge from image to video field, based on which Jin et al. [25] introduce a transformation block to adaptively transfer knowledge in the pre-trained CLIP from image-text to video-text. This adaptation is also achieved by a teacher-student framework in TeachClip [26]. Considering the feature redundancy issue, DRL [27] designs a loss in channel decorrelation regularization to improve the quality of video representation. However, these methods suffer from ill-suited cross-modal matching due to the information imbalance between videos and texts.
@@
To improve cross-modal matching, existing works, such as BIKE [11] and SHE-NET [12], enhance text expression with lexical and syntax analysis. Although Large language models (LLMs), such as ChatGPT and LLaMA [28], are additionally introduced in the text rewriting tasks [29, 30, 31], visual details remain deficient in the rewritten texts through these text-to-text literal restatements. Generating captions from videos as auxiliaries is a brighter approach for text enhancement. These methods form new pairs using videos and derived captions to promote cross-modal representation learning [15]. However, carefully designed prompts are necessary, and later works manually craft various query prompts, asking LLMs to describe empirically selective concepts, such as special objectives [16] or actions [17]. The hand-drafted lexicon and prompts severely limit the rewrite quality, motivating us to explore a self-learning caption enhancement method. To the best of our knowledge, this paper is a pioneer in investigating automatic caption enhancement for videos to facilitate the comprehension of videos and texts.
Figure 3: Illustration of the CSI module
III Method
@@
Unlike traditional text augmentation that rewrites captions on the text side, our method (ExCae) automatically derives captions from videos by the CSI module and adaptively selects appropriate expressions (ECS module) to encode video for video-text matching (see Fig. 2). Video and text embeddings would be aligned with standard cross-modal contrastive loss, as most existing SOTA methods [24, 3]. CSI and ECS are self-learning modules whose details are introduced as follows.
III-A Caption Self-improvement (CSI)
In the CSI module, a Captioner is included, which is an mLLM asked to generate a set of captions based on a query prompt to enhance video understanding. As intermediaries for video-text matching, the captions derived from videos are expected to be informative and text-oriented. To achieve this goal, CSI performs an iterative prompt refinement (Fig. 3), gradually updating the query prompts of the Captioner by repeatedly inquiring a Prompt Engineer. In each iteration, the current query prompt is evaluated by caption scoring, an approximate assessment of VTR performance, resulting in the update of the prompt candidate. The Prompt engineer then rewrites the current best prompt, whose output is used as the latest query prompt of the Captioner. This prompt update iteration would end when no new candidates are added. Pseudo code is provided in Algorithm 1.
@@ III-A Caption Self-improvement (CSI)
We hope that the adjustment of the query prompt will be positively correlated with the increase in VTR results throughout the iterative process. Directly evaluating VTR results as feedback for prompt refinement involves multimodal encoder optimization, where slight fluctuations in embeddings can significantly affect the direction of prompt optimization. Therefore, this paper designs a scoring mechanism that evaluates the quality of generated captions, implicitly reflecting the performance of VTR. Denote the query prompt for the Captioner at iteration $t$ as $p^{t}$. The generated caption list for video $v_{i}$, using $p^{t}$ is denoted as $\mathbf{c}_{i}^{t}=\{c_{i,1}^{t},c_{i,2}^{t},\ldots,c_{i,K}^{t}\}$, where $K$ is the number of captions per video. To balance semantic fidelity and diversity, $\mathbf{c}_{i}^{t}$ is evaluated from two perspectives:
  1. 1.
    Semantic Consistency: Ensures $\mathbf{c}_{i}^{t}$ semantically align with the original text $t_{i}$, mitigating the presence of cluttered and vacuous descriptions.
  2. 2.
    Caption Diversity: Encourages multi-angle descriptions in $\mathbf{c}_{i}^{t}$, avoiding redundancy to maintain the original intention of expression enhancement.
The overall score for iteration $t$ is computed as:
$Score^{t}=\mathbb{E}_{i,k}[sim(\psi(c^{t}_{i,k}),\psi(t_{i}))]+div(\psi(\mathbf{c}^{t}_{i}))$ (1)
where $k\in[1,K]$ is the caption index, and $\psi(\cdot)$ represents a semantic extractor. $sim(\cdot)$ calculates the cosine similarity on semantics. $div(\cdot)$ quantifies diversity within $\mathbf{c}^{t}_{i})$:
$div(\psi(\mathbf{c}^{t}_{i}))=\frac{1}{K(K-1)}\sum_{p\neq q}\left(1-sim\left(\psi(c_{i,p}^{t}),\psi(c_{i,q}^{t})\right)\right)$ (2)
The higher value of $Score^{t}$ indicates better alignment with the original text and greater caption diversity. If $Score^{t}$ surpasses the historical records, the prompt candidate is updated by $p^{t}$, and the best prompt used for prompt refinement is updated accordingly.
@@
Current prompt engineering normally encodes the query prompts and adjusts the prompt tokens/embeddings through gradient descent [20, 21]. Although this parametric fine-tuning strategy is convenient for optimization, the adjusted embeddings could hardly be precisely re-projected to the text space, losing linguistic meaning and interpretability. Consequently, their generalization in video-text-related applications is limited. Inspired by language rewrites [31], this paper introduces a Prompt engineer to rewrite the query prompt. Take the current best prompt $p^{t}$ as input; the Prompt engineer is asked to rewrite $p^{t}$ with reference to the corresponding video-text pairs $(x,t)$ involved in this iteration. The rewritten prompt then inquires the Captioner in the next iteration, as shown in Fig. 3.
Algorithm 1 Caption Self-improvement (CSI) algorithm
1:  Input: Video-text pairs $(x,t)$, Initial query prompt $p_{0}$, Maximum iteration $T$
2:  Output: $Best\_prompt$
3:  Initialization:
4:      $Best\_prompt\leftarrow p_{0};$
5:      $Best\_score\leftarrow-\infty;$
6:  for $t=1,2,\dots,T$ do
7:     $p_{t}\leftarrow PE(Best\_prompt)$;      // Refine prompt
8:     if $Best\_prompt==p_{t}$ then
9:        break;                           // No update
10:     end if
11:     $\mathbf{c}\leftarrow Captioner(x,p_{t})$;         // Generate Captions
12:     Calculate $Score^{t}$ using Equation (1);
13:     if $Score^{t}>Best\_score$ then
14:        $Best\_score\leftarrow Score^{t};$
15:        $Best\_prompt\leftarrow p_{t};$ ;
16:     end if
17:  end for
III-B Expertized Caption Selection (ECS)
With the optimized CSI module, video-text pairs are updated as $((v,\tilde{c}),t)$, where $(v,\tilde{c})$ represents the video data $v$ and its generated captions $\tilde{c}$ using the optimal prompt, and $t$ denotes the original text corresponding to $v$. As displayed in Fig. 2 (b), $(v,\tilde{c})$ and $t$ are the video-side and text-side input data, respectively. A straightforward approach for cross-modal representation learning is directly matching the encoded representations of $(v,\tilde{c})$ and $t$. However, the information dilation of video modality brings confounded knowledge while improving generalization, increasing the difficulty of cross-modal matching (see Fig. 1).
@@ III-B Expertized Caption Selection (ECS)
TABLE I: Comparison with SOTA methods on MSR-VTT dataset (underline marks the best ViT-based results)
@@
IV Experiment
IV-A Datasets
MSR-VTT [34] is a large-scale dataset containing 10,000 video clips. Each video is described in 20 English sentences. This paper uses 9,000 clips for training and 1,000 clips for testing, as in the setup of previous work [35].
MSVD [36] consists of 1,970 videos. Each video contains around 40 English captions. In total, around 80K captions are provided. The dataset is derived by 1,200/100/670 for training, validation and testing, respectively.
DiDeMo [37] contains over 10,000 videos downloaded from Flickr. Each video has a caption recording detailed information on camera movement, temporal transition indicators, and activities. We split the dataset into training, validation and test sets, each containing 8,395, 1,065 and 1,004 videos, respectively.
TABLE II: Comparison with SOTA methods on MSVD and DiDeMo datasets (underline marks the best ViT-based results)